# Model Inversion Attacks on Vision-Language Models: Do They Leak What They Learn?

### v1

@@ 3.5 Inversion Loss Design for VLMs
Cross-entropy-based. This loss is widely used in MI attacks (Zhang et al. 2020; Chen et al. 2021; Qiu et al. 2024) to optimize $w$ such that the reconstruction has the highest likelihood for the target class under the model $M$. For VLMs, we adapt the cross-entropy loss $\mathcal{L}_{CE}$ for each target token $y_{i}$ as follows:
@@
Logit-based. Prior work shows that using cross-entropy loss in MI can lead to gradient vanishing (Yuan et al. 2023) or sub-optimal results (Nguyen et al. 2023b). To address this, (Yuan et al. 2023; Nguyen et al. 2023b) propose optimizing losses directly over logits of a target class. We adopt two such logit-based losses for VLMs: the Max-Margin Loss $\mathcal{L}_{MML}$ (Yuan et al. 2023) and the Logit-Maximization Loss $\mathcal{L}_{LOM}$ (Nguyen et al. 2023b) for a target token $y_{i}$:
@@
Here, $l_{y_{i}}$ is the logit corresponding to the target token $y_{i}$, $\lambda$ is a hyperparameter, $f_{y_{i}}=M^{pen}(\mathbf{t},G(w),y_{<i}),y_{i})$ where $M^{pen}()$ denotes the function that extracts the penultimate layer representations for a given input, and $f_{reg}$ is a sample activation from the penultimate layer $M^{pen}()$ computed using public images from $\mathcal{D}_{pub}$. Following (Nguyen et al. 2023b), the distribution of $f_{reg}$ is estimated over 2000 input pairs $(\mathbf{t},\mathbf{x}_{pub})$, where $\mathbf{x}_{pub}\in\mathcal{D}_{pub}$.
$\mathcal{L}_{MML}$ maximizes the logit of the correct token $y_{i}$ while penalizing the highest incorrect logit to mitigate gradient vanishing. On the other hand, $\mathcal{L}_{LOM}$ also maximizes the correct token’s logit to avoid sub-optimality, while additionally penalizing deviations in the penultimate activations to prevent unbounded logits problem.
4 Experiments
In this section, we evaluate the effectiveness of our 4 proposed model inversion attacks on 3 VLMs (i.e., LLaVA-v1.6, Qwen2.5-VL and MiniGPT-v2), 3 private datasets, 2 public datasets with an extensive evaluation spanning 5 metrics including the human evaluation.
4.1 Experimental Setting
Dataset.
Following standard model inversion (MI) setups(Zhang et al. 2020; Chen et al. 2021; Struppek et al. 2022; An et al. 2022; Nguyen et al. 2023b; Yuan et al. 2023; Struppek, Hintersdorf, and Kersting 2024; Qiu et al. 2024; Ho et al. 2024; Koh et al. 2024), we use facial and fine-grained classification datasets to evaluate our approach. Specifically, we conduct experiments on three datasets: FaceScrub (Ng and Winkler 2014), CelebA (Liu et al. 2015), and Stanford Dogs (Dataset 2011). The FaceScrub dataset contains 106,836 images across 530 identities. For CelebA, we select the top 1,000 identities with the most samples from the full set of 10,177 identities. Stanford Dogs comprises images from 120 dog breeds, serving as a representative fine-grained visual dataset.
Public Dataset and Image Generator.
For facial image reconstruction, we use FFHQ (Karras, Laine, and Aila 2019) as the public dataset $\mathcal{D}_{pub}$ and a pre-trained StyleGAN2 (Karras et al. 2020) trained on FFHQ. Following conventional MI (Struppek et al. 2022), we optimize in the latent space $w$ of StyleGAN2 to recover images $x=G(w)$. For Stanford Dogs experiments, we adopt AFHQ-Dogs (Choi et al. 2020) as $\mathcal{D}_{pub}$ to train the dog image generator.
VLMs.
We fine-tune LLaVA-v1.6-7B (Liu et al. 2024), Qwen2.5VL-7B (Bai et al. 2025), and MiniGPT-v2 (Chen et al. 2023) using VQA-Facescrub, VQA-CelebA, and VQA-StanfordDogs.
Evaluation Metrics.
To assess the quality of the inversion results, we adopt five metrics:
  • Attack accuracy. We compute the attack accuracy using three frameworks as described below. We strictly follow the evaluation frameworks in their original works with the detailed setups provided in the Supp. Higher accuracy indicates a more effective inversion attack.
    • Attack accuracy evaluated by conventional evaluation framework $\mathcal{F}_{DNN}$ ($AttAcc_{D}\uparrow$) (Zhang et al. 2020; Chen et al. 2021; Struppek et al. 2022; Nguyen et al. 2023b; Qiu et al. 2024). This is a conventional framework, where the evaluation models are standard DNNs trained on private dataset. Following (Struppek et al. 2022; Struppek, Hintersdorf, and Kersting 2024), we use InceptionNet-v3 (Szegedy et al. 2016) as the evaluation model to classify reconstructed images, and compute the $Top1$ and $Top5$ based on whether the predicted label match the target label.
    • Attack accuracy evaluated by MLLM-based evaluation framework $\mathcal{F}_{MLLM}$ ($AttAcc_{M}\uparrow$). (Ho et al. 2025) demonstrate that $\mathcal{F}_{MLLM}$ can achieve better alignment with human evaluation. Unlike the conventional framework $\mathcal{F}_{DNN}$, which relies on the classification predictions of standard DNNs trained on private datasets, this metric leverages powerful MLLMs to evaluate the success of MI-reconstructed by referencing the corresponding private images.
    • Attack accuracy evaluated by human $\mathcal{F}_{Human}(AttAcc_{H}\uparrow)$. Following existing studies (An et al. 2022; Nguyen et al. 2023b), we conduct the user study on Amazon Mechanical Turk. Participants are asked to evaluate the success of MI-reconstructed by referencing the corresponding private images (Details in the Supp).
  • Feature distance. We compute the $l_{2}$ distance between the feature representations of the reconstructed and the private training images (Struppek et al. 2022). Lower values indicate higher similarity and better inversion quality.
    • $\delta_{face}$. Features are extracted by a pre-trained FaceNet model (Schroff, Kalenichenko, and Philbin 2015).
4.2 Results
$\mathcal{L}_{inv}$ $tAcc_{M}\uparrow AttA$$tAcc_{D}\uparrow$ $\delta_{face}\downarrow$ $\delta_{eval}\downarrow$ $Top1$ $Top5$ TMI $\mathcal{L}_{CE}$ 37.78% 17.71% 39.79% 0.8939 147.35 $\mathcal{L}_{MML}$ 39.98% 17.31% 38.51% 0.9065 193.14 $\mathcal{L}_{LOM}$ 44.34% 21.77% 44.69% 0.8488 141.87 TMI-C $\mathcal{L}_{CE}$ 21.77% 6.39% 18.58% 1.0911 636.50 $\mathcal{L}_{MML}$ 25.99% 6.51% 18.82% 1.0659 205.71 $\mathcal{L}_{LOM}$ 31.16% 9.32% 24.22% 1.0221 457.49 SMI $\mathcal{L}_{CE}$ 40.97% 18.25% 41.11% 0.8682 144.53 $\mathcal{L}_{MML}$ 55.52% 32.83% 60.12% 0.7569 137.43 $\mathcal{L}_{LOM}$ 59.17% 33.47% 61.89% 0.7465 140.83 SMI-AW $\mathcal{L}_{CE}$ 44.17% 22.33% 46.63% 0.8464 145.29 $\mathcal{L}_{MML}$ 57.15% 34.91% 61.84% 0.7444 138.24 $\mathcal{L}_{LOM}$ 59.25% 36.98% 64.69% 0.7286 135.90
Table 1: Comparison of performance metrics across four inversion strategies using LLaVa-1.6-7B fine-tuned on the FaceScrub dataset, evaluated with three identity losses.
We report attack results on the FaceScrub dataset in Table 1, evaluating four MI strategies under three inversion losses using LLaVa-1.6-7B. The results show that sequence-based mode inversion methods consistently outperform token-level MI approaches across all evaluation metrics. Among them, SMI-AW, when combined with the $\mathcal{L}_{LOM}$, achieves the highest performance. This highlights the advantage of employing adaptive token-wise weights that are dynamically updated at each inversion step. Using this method, we achieve an attack accuracy of 59.25% under $\mathcal{F}_{MLLM}$ while other distance metrics such as $\delta{face}$ and $\delta_{eval}$ are the lowest (where lower is better).
Results on additional datasets, including CelebA and Stanford Dogs, are shown in Table 2 and Table 3, using the logit maximization loss. We achieve high attack success rates, with attack accuracies of 66.91% on CelebA and 77.40% on Stanford Dogs. These findings are consistent with results on the FaceScrub dataset, where SMI-AW consistently achieves the highest attack performance across all evaluation metrics.
We further evaluate our proposed method on Qwen2.5-VL-7B and MiniGPT-v2, using the FaceScrub dataset. The results reinforce the generalizability of our findings, demonstrating that VLMs are broadly vulnerable to model inversion attacks. These results underscore the severity of this vulnerability and raise a significant alarm about the susceptibility of VLMs to inversion-based privacy breaches.
Method $tAcc_{M}\uparrow AttA$$tAcc_{D}\uparrow$ $\delta_{face}\downarrow$ $\delta_{eval}\downarrow$ $Top1$ $Top5$ TMI 39.74% 15.31% 33.14% 1.0195 428.66 TMI-C 18.73% 3.63% 10.29% 1.2370 446.90 SMI 64.93% 38.30% 63.69% 0.8294 416.34 SMI-AW 66.91% 40.83% 65.84% 0.8133 415.25
Table 2: We report the results on the CelebA dataset across four inversion strategies with $\mathcal{L}_{LOM}$.
Method $tAcc_{M}\uparrow AttA$$tAcc_{D}\uparrow$ $\delta_{eval}\downarrow$ $Top1$ $Top5$ TMI 61.46% 40.31% 70.21% 102.40 TMI-C 48.54% 29.69% 59.79% 102.23 SMI 75.94% 53.65% 82.19% 76.98 SMI-AW 77.40% 58.33% 86.04% 78.61
Table 3: We report the results on the Stanford Dogs dataset across four inversion strategies with $\mathcal{L}_{LOM}$.
$M$ $tAcc_{M}\uparrow AttA$$tAcc_{D}\uparrow$ $\delta_{face}\downarrow$ $\delta_{eval}\downarrow$ $Top1$ $Top5$ MiniGPT-v2 50.80% 15.26% 34.69% 0.8909 161.35 Qwen2.5-VL 36.42% 14.91% 31.37% 1.0115 144.92
Table 4: We report the results of Qwen2.5-VL-7B and MiniGPT-v2 on the Facescub dataset. Here we use SMI-AW with the logit-maximization loss $\mathcal{L}_{LOM}$.
4.3 Analysis
To better understand why token-based MI methods underperform compared to sequence-based approaches, we analyze the match rate between the final reconstructed images $M(\mathbf{t},G(w^{*}))$ and the corresponding target textual answers $y$. Specifically, we define the match rate as the percentage of reconstructed images for which the target answer $y$ appears as a substring of the predicted text associated with the image. In other words, it reflects the proportion of reconstructions whose generated text aligns with the target textual answer at the end of the inversion process.
The results, shown in Figure 2, reveal a clear distinction between the two types of methods. Token-based MIs exhibit poor convergence behavior, with match rates ranging from 60% to 79% for TMI, and dropping below 30% for TMI-C. In contrast, sequence-based methods such as SMI and SMI-AW achieve match rates exceeding 95%, indicating more reliable alignment between reconstructed images and their textual targets. It is important to note that a high match rate does not necessarily imply a successful attack, as the optimization may overfit or converge to a poor local minimum. Nevertheless, a higher match rate generally correlates with a greater likelihood of a successful identity inversion attack.
4.4 Qualitative Results
Figure 3 shows qualitative results demonstrating the effectiveness of our method. Using SMI-AW with $\mathcal{L}_{LOM}$, the reconstructed images from the LLaVA-v1.6-7B model (second row) closely resemble the corresponding identities in $\mathcal{D}{priv}$ (first row). This strong visual similarity highlights the ability of our model inversion approach to recover identifiable features from the training data.
Figure 3: Qualitative results on the Facescrub dataset using the SMI-AW and $\mathcal{L}_{LOM}$. The first row shows images from the private training dataset, while the second row presents the reconstructed images corresponding to each individual in the first row. The visual similarity between the original and reconstructed images demonstrates the effectiveness of our inversion method in recovering private training data. More reconstructed images can be found in Supp.
4.5 Human Evaluation
We further conduct human evaluation on reconstructed images using two datasets Facescrub and CelebA. Each user study involves 4,240 participants for the FaceScrub dataset and 8,000 participants for the CelebA dataset. The results show that 57.74% to 75.31% of the reconstructed samples are deemed successful attacks, i.e., human annotators recognize the generated images as depicting the same identity as those in the private image set (see Table 5). This highlights the alarming potential of such inversion attacks to compromise sensitive identity information.
VLM $\mathcal{D}_{priv}$ $AttAcc_{H}\uparrow$
LLaVA-v1.6-7B Facescrub 75.31%
MiniGPT-v2 61.84%
Qwen2.5-VL 57.74%
LLaVA-v1.6-7B CelebA 61.95%
Table 5: Human evaluation results. We evaluate our method using SMI-AW with $\mathcal{L}_{LOM}$, the private datasets $\mathcal{D}_{priv}$ are FaceScrub and CelebA.
4.6 Evaluation with Publicly Released VLM
In the experiments above, we fine-tuned the target model using a private training dataset following prior MI work on conventional DNNs. In this section, we extend our analysis to the publicly available LLaVA-v1.6-7B model, aiming to reconstruct potential training images directly from it.
Figure 4 shows the results of our best setup of MI attack, SMI-AW using the logit maximization loss. The target is to reconstruct images of Donald Trump that appear in the training dataset of the LLaVA-v1.6-7B model. We visualize two image pairs: in each pair, the left image is the reconstruction generated from the publicly available model, and the right image is a training image of Donald Trump. The visual similarity between the pairs indicates that the pre-trained VLM may reveal identifiable information from its training data, exposing a vulnerability to model inversion attacks.
Figure 4: We reconstruct images of Donald Trump from the pre-trained LLaVA-v1.6-7B model. We use SMI-AW with $\mathcal{L}_{LOM}$ to reconstruct images. For each pair, the right image shows a training image of Donald Trump, while the left image presents the reconstruction obtained via our model inversion attack. This result illustrates that the pre-trained VLM is vulnerable to training data leakage through model inversion. More results can be found in Supp.
5 Conclusion
This study pioneers the investigation of model inversion attacks on Vision-Language Models, demonstrating for the first time their susceptibility to leaking private visual training data. Our novel token-based and sequence-based inversion strategies reveal significant privacy risks across state-of-the-art and publicly available VLMs. These findings underscore the urgent need for robust privacy safeguards as VLMs become more prevalent in real-world applications.
Supplementary material
1 Research Reproducibility Details
1.1 Hyperparameters
To fine-tune the VLMs, we follow the standard hyperparameters provided in the official implementations of LLaVA-v1.6-Vicuna-7B [^1] https://github.com/haotian-liu/LLaVA (Liu et al. 2024), Qwen2.5-VL-7B [^2] https://github.com/QwenLM/Qwen2.5-VL (Bai et al. 2025), and MiniGPT-v2333https://github.com/Vision-CAIR/MiniGPT-4 (Chen et al. 2023). Fine-tuning is conducted on the VQA-FaceScrub, VQA-CelebA, and VQA-StanfordDogs datasets.
For the attacks, we use $N=70$ inversion steps for experiments on the LLaVA-v1.6-7B model, and $N=100$ for MiniGPT-v2 and Qwen2.5-VL-7B. The inversion update rate $\beta=0.05$. We set the confidence threshold $p_{thres}=0.999$ for all experiments using the logit maximization loss $\mathcal{L}_{LOM}$. Additional results with varying values of $p_{thres}$ are provided in the ablation study section (Supp).
To compute the regularization term $f_{reg}$ in Eqn. (8), we follow (Nguyen et al. 2023b) by using 2,000 images from a public dataset $\mathcal{D}_{pub}$ to estimate the mean and variance of the penultimate layer activations of the VLMs.
1.2 Computational Resources
All experiments were conducted on NVIDIA RTX A6000 Ada GPUs running Ubuntu 20.04.2 LTS, equipped with AMD Ryzen Threadripper PRO 5975WX 32-core processors. The environment setup for each model is provided in the official implementations of the VLMs, including: LLaVA-v1.6-Vicuna-7B (Liu et al. 2024), Qwen2.5-VL-7B (Bai et al. 2025), and MiniGPT-v2 (Chen et al. 2023).
To evaluate $AttAcc_{M}$, we strictly follow the protocol in (Ho et al. 2025), using the Gemini 2.0 Flash API. In total, we evaluate 95,200 MI-reconstructed images for our main experiments (main paper).
2 Additional results
2.1 Extended Evaluation on Publicly Released VLM
(a) Beyoncé
(b) Ed Sheeran
(c) Kristen Stewart
(d) Barack Obama
In this section, we extend our analysis to the publicly available LLaVA-v1.6-7B model (Liu et al. 2024), aiming to reconstruct training images from accessing the model only.
Figure S.1 shows the results of our best setup of MI attack, SMI-AW using the logit maximization loss $\mathcal{L}_{LOM}$. The target is to reconstruct images of celebrities that appear in the training dataset of the LLaVA-v1.6-7B model. To reconstructed images from the model, we use the textual input $t$ = “Identify the person in the image and return only their name?” and the target textual answer is a celebrity’s name, i.e $y$ = “Beyoncé”.
2.2 Additional Qualitative Results
Reconstructed images from the FaceScrub dataset using three VLMs, LLaVA-v1.6-7B, MiniGPT-v2, and Qwen2.5-VL, are shown in Figure S.2, Figure S.3, and Figure S.4, respectively. For the CelebA and Stanford Dogs datasets, reconstructed images using LLaVA-v1.6-7B are presented in Figure S.5 and Figure S.6. All reconstructions are generated using SMI-AW with the logit maximization loss $\mathcal{L}_{LOM}$.
Figure S.2: Qualitative results on the Facescrub dataset using the SMI-AW and $\mathcal{L}_{LOM}$, $M$ = LLaVA-v1.6-7B. For each pair, the left column shows images from the private training dataset, while the right column presents the reconstructed images corresponding to each individual in the left column.
Figure S.3: Qualitative results on the Facescrub dataset using the SMI-AW and $\mathcal{L}_{LOM}$, $M$ = MiniGPT-v2. For each pair, the left column shows images from the private training dataset, while the right column presents the reconstructed images corresponding to each individual in the left column.
Figure S.4: Qualitative results on the Facescrub dataset using the SMI-AW and $\mathcal{L}_{LOM}$, $M$ = Qwen2.5-VL. For each pair, the left column shows images from the private training dataset, while the right column presents the reconstructed images corresponding to each individual in the left column.
Figure S.5: Qualitative results on the CelebA dataset using the SMI-AW and $\mathcal{L}_{LOM}$, $M$ = LLaVA-v1.6-7B. For each pair, the left column shows images from the private training dataset, while the right column presents the reconstructed images corresponding to each individual in the left column.
3 Ablation Study
3.1 Ablation Study on $p_{thres}$
We conduct an ablation study to investigate the effect of setting $p_{thres}$ in SMI-AW. Here, we use $M$ = LLaVA-v1.6-7B, $\mathcal{D}_{priv}$ = Facescrub, $\mathcal{L}_{inv}=\mathcal{L}_{LOM}$. As shown in Table S.1, using a higher threshold to focus on tokens with low confidence scores consistently improves attack performance across all evaluation metrics. For all experiments in main paper, we use $p=0.999$.
$p_{thres}$ $tAcc_{M}\uparrow AttA$$tAcc_{D}\uparrow$ $\delta_{face}\downarrow$ $\delta_{eval}\downarrow$ $Top1$ $Top5$ 0.999 59.83% 37.17% 65.31% 0.7349 135.81 0.98 57.05% 34.32% 61.23% 0.7486 137.27 0.95 56.96% 33.21% 61.86% 0.7495 136.71
Table S.1: Ablation study on $p_{thres}$ for adaptive token weights in SMI-AW. Here, we use $M$ = LLaVA-v1.6-7B, $\mathcal{D}_{priv}$ = Facescrub, $\mathcal{L}_{inv}=\mathcal{L}_{LOM}$. Using a higher threshold to focus on tokens with low confidence scores consistently improves attack performance across all evaluation metrics.
3.2 Error Bar
We repeat each experiment three times using different random seeds and report the results in Table S.2. Specifically, we use $M$ = LLaVA-v1.6-7B, $\mathcal{D}_{priv}$ = Facescrub, and $p_{thres}=0.999$. The results demonstrate that our attacks have low standard deviation.
$Top1$ $Top5$
SMI-AW 59.39 $\pm$ 0.39% 37.00 $\pm$ 0.17% 65.01 $\pm$ 0.31% 0.7318 $\pm$ 0.0031 135.84 $\pm$ 0.05
Table S.2: Error bars for our two model inversion strategies SMI and SMI-AW. Each experiment was repeated 3 times, and we report the mean and standard deviation of the attack performance. Here, we use $M=$ LLaVa-v1.6-7B, $\mathcal{D}_{priv}=$ Facescrub, $p_{thres}=0.999$. All inversion strategies are combined with logit maximization loss $\mathbf{L}_{LOM}$.
4 Experimental setting
@@ 4.1 Evaluation metrics
    • Attack accuracy evaluated by conventional evaluation framework $\mathcal{F}_{DNN}$ ($AttAcc_{D}\uparrow$) (Zhang et al. 2020; Chen et al. 2021; Struppek et al. 2022; Nguyen et al. 2023b; Qiu et al. 2024). Following (Struppek et al. 2022; Struppek, Hintersdorf, and Kersting 2024), we use InceptionNet-v3 (Szegedy et al. 2016) as the evaluation model. For a fair comparison, we use the identical checkpoints of InceptionNet-v3 for Facescrubs, CelebA and Stanford Dogs from (Struppek et al. 2022) for evaluation of each dataset. We report Top-1 and Top-5 Accuracy.
    • Attack accuracy evaluated by MLLM-based evaluation framework $\mathcal{F}_{MLLM}$ ($AttAcc_{M}\uparrow$). (Ho et al. 2025) demonstrate that $\mathcal{F}_{MLLM}$ can achieve better alignment with human evaluation than $\mathcal{F}_{DNN}$ ($AttAcc_{D}\uparrow$ by mitigating Type-I adversarial transferability. The evaluation involves presenting a reconstructed image (image A) and a set of private reference images (set B) to an MLLM (e.g., Gemini 2.0 Flash), and prompting it with the question: “Does image A depict the same individual as images in set B?” If the model responds “Yes”, the attack is considered successful. An example query is shown in Fig. S.7.
    • Attack accuracy evaluated by human $\mathcal{F}_{Human}(AttAcc_{H}\uparrow)$. Following existing studies (An et al. 2022; Nguyen et al. 2023b), we conduct the user study on Amazon Mechanical Turk. Participants are asked to evaluate the success of MI-reconstructed by referencing the corresponding private images. Similar to $\mathcal{F}_{MLLM}$, it involves presenting an image A and a set of images B. They are asked to answer “Yes” or “No” to indicate whether image A depicts the same identity as images in set B (see Fig. S.7). Each image pair is shown in a randomized order and displayed for up to 60 seconds. Each user study involves 4,240 participants for the FaceScrub dataset and 8,000 participants for the CelebA dataset.
  • Feature distance. We compute the $l_{2}$ distance between the feature representations of the reconstructed and the private training images (Struppek et al. 2022). Lower values indicate higher similarity and better inversion quality.
@@
    • $\delta_{face}$. Features are extracted by a pre-trained FaceNet model (Schroff, Kalenichenko, and Philbin 2015).

### v2

@@ 4.1 Inversion Loss Design for VLMs
Cross-entropy-based. This loss is widely used in MI attacks [zhang2020secret, chen2021knowledge, qiu2024closer] to optimize $w$ such that the reconstruction has the highest likelihood for the target class under the model $M$. For VLMs, we adapt the cross-entropy loss $\mathcal{L}_{CE}$ for each target token $y_{i}$ as follows:
@@
Logit-based. Prior work shows that using cross-entropy loss in MI can lead to gradient vanishing [yuan2023pseudo] or sub-optimal results [nguyen2023re]. To address this, yuan2023pseudo and nguyen2023re propose optimizing losses directly over logits of a target class. We adopt two such logit-based losses for VLMs: the Max-Margin Loss $\mathcal{L}_{MML}$ [yuan2023pseudo] and the Logit-Maximization Loss $\mathcal{L}_{LOM}$ [nguyen2023re] for a target token $y_{i}$:
@@
Here, $l_{y_{i}}$ is the logit corresponding to the target token $y_{i}$, $\lambda$ is a hyperparameter, $f_{y_{i}}=M^{pen}(\mathbf{t},G(w),y_{<i})$ where $M^{pen}()$ denotes the function that extracts the penultimate layer representations for a given input, and $f_{reg}$ is a sample activation from the penultimate layer $M^{pen}()$ computed using public images from $\mathcal{D}_{pub}$. Following [nguyen2023re], the distribution of $f_{reg}$ is estimated over 2000 input pairs $(\mathbf{t},\mathbf{x}_{pub})$, where $\mathbf{x}_{pub}\in\mathcal{D}_{pub}$. $\mathcal{L}_{MML}$ maximizes the logit of the correct token $y_{i}$ while penalizing the highest incorrect logit to mitigate gradient vanishing. On the other hand, $\mathcal{L}_{LOM}$ also maximizes the correct token’s logit to avoid sub-optimality, while additionally penalizing deviations in the penultimate activations to prevent unbounded logits problem.
@@ 4.2 Evaluation metrics
    • Attack accuracy evaluated by conventional evaluation framework $\mathcal{F}_{DNN}$ ($AttAcc_{D}\uparrow$) [zhang2020secret, chen2021knowledge, struppek2022plug, nguyen2023re, qiu2024closer]. Following [struppek2022plug, struppekcareful], we use InceptionNet-v3 [szegedy2016rethinking] as the evaluation model. For a fair comparison, we use the identical checkpoints of InceptionNet-v3 for Facescrubs, CelebA and Stanford Dogs from [struppek2022plug] for evaluation of each dataset. We report Top-1 and Top-5 Accuracy.
    • Attack accuracy evaluated by MLLM-based evaluation framework $\mathcal{F}_{MLLM}$ ($AttAcc_{M}\uparrow$). [ho2025revisitingmodelinversionevaluation] demonstrate that $\mathcal{F}_{MLLM}$ can achieve better alignment with human evaluation than $\mathcal{F}_{DNN}$ ($AttAcc_{D}\uparrow$ by mitigating Type-I adversarial transferability. The evaluation involves presenting a reconstructed image (image A) and a set of private reference images (set B) to an MLLM (e.g., Gemini 2.0 Flash), and prompting it with the question: “Does image A depict the same individual as images in set B?” If the model responds “Yes”, the attack is considered successful. An example query is shown in Fig. S.13.
    • Attack accuracy evaluated by human $\mathcal{F}_{Human}(AttAcc_{H}\uparrow)$. Following existing studies [an2022mirror, nguyen2023re], we conduct the user study on Amazon Mechanical Turk. Participants are asked to evaluate the success of MI-reconstructed by referencing the corresponding private images. Similar to $\mathcal{F}_{MLLM}$, it involves presenting an image A and a set of images B. They are asked to answer “Yes” or “No” to indicate whether image A depicts the same identity as images in set B (see Fig. S.13). Each image pair is shown in a randomized order and displayed for up to 60 seconds. Each user study involves 4,240 participants for the FaceScrub dataset and 8,000 participants for the CelebA dataset.
  • Feature distance. We compute the $l_{2}$ distance between the feature representations of the reconstructed and the private training images [struppek2022plug]. Lower values indicate higher similarity and better inversion quality.
@@
    • $\delta_{face}$. Features are extracted by a pre-trained FaceNet model [schroff2015facenet].